# Combining Static and Dynamic Approaches for Mining and Testing Constraints for RESTful API Testing

### v1

@@ Abstract.
In API testing, deriving logical constraints on API response bodies is crucial in generating the test cases to cover various aspects of RESTful APIs. However, existing approaches are limited to dynamic analysis in which constraints are extracted from the execution of APIs as part of the system under test. The key limitation of such a dynamic approach is its under-estimation in which inputs in API executions are not sufficiently diverse to uncover actual constraints on API response bodies. In this paper, we propose to combine a novel static analysis approach (in which the constraints for API response bodies are mined from API specifications), with the dynamic approach (which relies on API execution data). We leverage large language models (LLMs) to comprehend the API specifications, mine constraints for response bodies, and generate test cases. To reduce LLMs’ hallucination, we apply an Observation-Confirmation (OC) scheme which uses initial prompts to contextualize constraints. Our empirical results show that LLMs with OC prompting achieve high precision in constraint mining with the average of 91.2%. When combining static and dynamic analysis, our tool, RBCTest, achieves a precision of 78.5%. RBCTest detects 107 constraints that the dynamic approach misses and 46 more precise constraints. We also use its generated test cases to detect 21 mismatches between the API specification and actual response data for 8 real-world APIs. Four of the mismatches were, in fact, reported in developers’ forums.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; Mar 15, 2025; XXX, XXXcopyright: none
@@ 1. Introduction
By adhering to the principles of Representational State Transfer (REST), the RESTful APIs provide a standardized way for interoperability among components and software systems. RESTful API testing helps identify and resolve several issues, ensuring that APIs perform as expected (Viglianisi et al., 2020; Martin-Lopez et al., 2021b; Arcuri, 2019; Liu et al., 2022; Alonso et al., 2022; Atlidakis et al., 2019). It also helps verify that APIs adhere to specifications and handle edge cases gracefully. Among techniques for API testing, black-box testing uses the OpenAPI Specification (OAS) as a basis to generate test cases and data (Le et al., 2024; Karlsson et al., 2020; Martin-Lopez et al., 2021b). The state-of-the-art API testing approaches are focused on status code (Viglianisi et al., 2020; Le et al., 2024; kat, 2024; pos, 2024) and schema validation (Viglianisi et al., 2020; kat, 2024; pos, 2024), even with rule extraction using human-readable descriptions in the OAS (Kim et al., 2023). In status code validation, each HTTP request returns a response with a status code, a three-digit integer, indicating the outcome of the HTTP request. Current testing approaches define an oracle for a test case by validating whether the response status code matches the expected value. In contrast, schema validation ensures the correctness of the response data by checking it against a specified schema. This involves verifying the presence of all required properties and ensuring the data types of these properties match their schema in the specification.
@@
While status code and schema validation effectively cover aspects of data representation and status checking, they may overlook the logical correctness and validity of the response data from the APIs, which is essential for software reliability. For example, an API request for a customer older than 18 receiving a response for one younger than 18 would not be detected by just validating the status code or schema. Deriving logical constraints on API response bodies is essential for generating test cases to cover RESTful APIs.
@@
The state-of-the-art approaches for mining constraints on API response bodies focus only on dynamic analysis in which the constraints are extracted from the execution data of the system under test (SUT). AGORA (Alonso et al., 2023) automatically detects invariants—properties that should consistently hold true. To identify invariants, which serve as logical constraints on API response bodies, it extends Daikon (Ernst et al., 2007), a dynamic instrumenter used to detect invariants during execution. As with dynamic analysis, it under-estimates the constraints due to the lack of diverse inputs to cover different aspects of API response bodies. For example, the inputs of the APIs might not be diverse enough to discover that the minimum age for an operation on a website is 18. Another limitation is that it requires the SUT operate accurately to extract the constraints from inputs and outputs.
We propose RBCTest, a combined approach between a novel LLM-based static approach to mine the constraints of API response bodies from the API specification, and the dynamic approach as in AGORA (Alonso et al., 2023). Both approaches complement to each other in the following ways. First, constraint mining remains feasible even if only the API specification or execution data is available. For instance, when an API specification exists but the APIs and the system utilizing them are still under testing and development and may not function correctly. Conversely, in a regression testing scenario without up-to-date specification, the SUT with APIs from previous versions has been functioning, while the current version is still under development. In such cases, execution data from the previous version can be used to extract constraints to generate test cases for regression testing. Second, the specification often provides API details, including request bodies sent to the API and response bodies returned for each operation. This enables mining constraints on the API’s response bodies to uncover more comprehensive information about these constraints, which are overlooked by the dynamic approach. In contrast, due to actual execution, the runtime data helps derive more detailed constraints not defined in the specifications.
@@
To mine constraints, we leverage the ability of large language models (LLMs) to comprehend natural language descriptions in API specifications. Constraints on API response bodies are inferred from different sources, including response properties, response schema, operations, and request parameters. We also harness LLMs’ proficiency to create test cases from the mined constraints to verify if the SUT correctly returns the content satisfying the mined constraints. Moreover, we apply an Observation-Confirmation scheme. We divide the task of mining constraints into two phases: observation and confirmation. The initial prompt contextualizes the description of constraints, enabling the next prompt to more accurately decide their presence. As another issue in LLMs’ exploration capability, they could produce resulting constraints that are not true. Thus, we enhance RBCTest with two extra mechanisms. First, before requesting the LLM to make observations concerning constraints on parameters, we perform a filtering process to keep only the valid ones. Second, after generating test cases for mined constraints, we add a semantic verifier to verify those test cases against the examples specified in the OAS file. The idea is that such examples tend to be correct because they illustrate the descriptions on the data types or the data. For example, the OAS could give ”March” as a valid month. If such a correct example does not pass a test case generated by RBCTest based on the mined constraint(s), the test case must be incorrect, which is caused by incorrect constraint(s). Thus, our verifier will discard them, leading to an improved precision.
@@
We evaluated RBCTest on two datasets, one from the baseline (Alonso et al., 2023) (AGORA dataset) with 11 operations in 7 API services and our RBCTest dataset collected from 8 real-world API services consisting of 59 endpoints and 83 operations (web, 2025). Our empirical results show that GPT-4-turbo and our OC prompting achieves high precision in constraint mining with the average of 91.2%. RBCTest achieves a precision of 78.5%, and detects 107 constraints that the dynamic approach misses and 46 more precise constraints. We also leverage its generated tests to detect the mismatches between the API specifications and actual execution of the SUTs. A detected mismatch indicates a fault in the SUT or that the specification does not reflect the SUT. We report 21 actual faults found in 8 real-world applications, including 4 issues reported by users on GitLab Forum (git, 2024b, c, d, a). In brief, this paper makes the following contributions:
@@
1. RBCTest: [A combination of static and dynamic approaches] for constraint mining and test generation for API response bodies by using API specifications.
@@
2. [A manually-verified benchmark] for API response bodies is available (web, 2025) for future research on API testing approaches.
3. [An extensive evaluation] showing RBCTest outperforming the state-of-the-art baselines.
2. Motivating Example
2.1. Example and Observations
@@
To illustrate the challenges and motivate our approach, we use Stripe, an online payment service, streamlining the process of charging customers via APIs. Figure 1(b) shows a simplified description from the API specification (OAS), detailing the GET operation for retrieving past charges. This API enables users to retrieve charging records within a specific time interval (lines 11-12). The time interval is defined by the gt (greater than) and lt (less than) parameters, representing the lower and upper bounds of the time range, respectively. Users can also specify the customer for whom they wish to retrieve charging history (lines 13-16). A successful request returns a response with a status code of 200 and a response body with data. The structure of the response data is outlined in the schema in Figure 1(a), which consists of a list of charge objects, each associated with various properties, e.g., amount, created timestamp, currency, and others. For instance, a GET request to the /v1/charges endpoint with parameters like ‘created[gt]=1679090500&customer=cus_idA’ returns a list of charges that satisfy the given conditions. An example response object is displayed in Figure 2. The content of the response is constrained by the actual input parameters in the request. Thus, a complete testing process needs to verify the response content in addition to the returned status code. For example, a test case can check if the created field of each returned charge object falls within the specified interval and ensuring that the customer field matches the requested ID.
Observation 1 (Constraints from input parameters).
The response data is constrained by the input parameters from the request. When testing, in addition to verifying the status code, testers need to verify the response data.
In Figure 1(a), the descriptions on the properties of the returned charge objects define certain constraints on the attributes of those objects. For example, the amount property must be a positive integer, with a maximum value of eight digits. The currency attribute has a three-letter lowercase code (e.g., usd).
Observation 2 (Constraints within response body).
Natural language descriptions on operations express logical constraints on operations, their responses and others, formatting requirements, or value range limitations that must be validated during API testing.
@@
2.2. State-of-the-art Approaches
In API test case generation, validation typically falls into two main categories: status code validation and schema validation.
@@ 2.2. State-of-the-art Approaches
Status Code Validation: Each HTTP request is returned with a response containing a status code and data. The status code, a 3-digit integer, indicates the outcome of the HTTP request. 2xx codes signify a successful request. Conversely, 4xx codes indicate errors, such as a bad syntax request or invalid input values. For instance, in testing the API ’GET/user_information’ with various valid and invalid user IDs, testers would expect the API to return status codes 200 or 404 based on the inputs provided. This validation method is widely used in automation testing tools, e.g., Postman (pos, 2024) and Katalon (kat, 2024), RestTestGen (Viglianisi et al., 2020), or KAT (Le et al., 2024).
@@
Schema Validation: Schema validation ensures the response correctness by checking it against a predefined response schema. This verifies the presence of all required properties and the consistency of property data types with their specifications. A lack of required data or a mismatch in data types indicates errors in API services. Tools, e.g., RestTestGen (Viglianisi et al., 2020), leverage external libraries, such as ‘swagger-schema-validator’, to facilitate schema validation.
@@
While status code and schema validation effectively cover aspects of data representation and status checking, they may overlook the logical correctness and validity in the response data. For instance, if an API request for a charge in 2025 returns one from 2024, or if the charge amount is negative, these issues would not be detected by merely validating the status code or schema.
To address that, the state-of-the-art dynamic approach, AGORA (Alonso et al., 2023) extends Daikon (Ernst et al., 2007), a dynamic instrumenter, to infer the invariants from the values extracted from the execution of the SUT using the APIs. AGORA considers these derived invariants as the constraints for the response bodies. However, inherent from the nature of a dynamic approach, the quality of the derived invariants depends on the values observed during the execution, which might not be diverse enough to reveal the correct constraints. For example, for all the inputs, the charge values might never reach 99,999,999, thus, Daikon returns the maximum charge that is less than that value.
2.3. Key Ideas
@@
From the above observations, we draw the following key ideas to design RBCTest, an approach for mining the constraints of response bodies and then generating test cases to validate them.
@@
Figure 3. Static Constraint Mining with LLMs
@@
Key Idea 1 [Combining Static and Dynamic Approaches to Mine Constraints of API’s response bodies]. The first component of RBCTest is a novel LLM-based static approach to mine the constraints for API’s response bodies from the OpenAPI Specification (OAS). The constraints on the response data can be found in the specification in either the descriptions of the operations (e.g., line 5 of Figure 1(a)) or the descriptions of the schema for the response data (lines 2–24 in Figure 1(b)). For example, the description in Figure 1 states that ‘the charges are returned in a sorted order with the most recent ones appearing first’. The schema for the returned values in Figure 1 also provides us several constraints on the parameters (lines 5–6) as well as the description of the returned object as a whole (line 2). For example, the amount of a charge is a positive number with up to 8 digits and such value must be of integer.
We integrate AGORA as the dynamic component, leveraging execution data from the SUT to mine invariants for API response bodies. The static and dynamic components complement each other. First, constraint mining remains feasible even if only the API specification or execution data is available. Second, the static component facilitates mining constraints on API response bodies from API specifications, capturing broader constraint information that the dynamic approach might miss. Conversely, runtime data, derived from actual executions, can help refine constraints, providing more precise insights than those typically found in API specifications.
@@
Key Idea 4 [Filter and Semantic Verifier]. Before asking the LLM to analyze constraints on parameters, we first perform a filtering process to remove invalid constraints. After generating test cases based on the mined constraints, we then introduce a semantic verifier to check these test cases against the examples in the OAS. The rationale is that these examples are typically accurate, as they represent the valid data types or values. For instance, an example of \$999,999.99 is used to illustrate a positive number for amount. We can validate the generated test cases against such examples. If a test case fails to validate a given correct example, it suggests that the mined constraint may be incorrect.
3. Static Constraint Mining
To mine the constraints, we observe that the specification for an API endpoint includes two main parts: the request specification (the input of the API) and response schema specification (the output of the API). (1) A request specification determines how to call an API endpoint, detailing the required input parameters, their roles, and the parameters within the request body along with their respective roles. It might also contain the description of the API operations. (2) The response schema specification provides instructions on the response data, including each property in the response data, its description, datatype, nullability, and other relevant details. These parts contain descriptions that are the targets for constraint mining.

### v3

# RBCTest: Leveraging LLMs to Mine and Verify Oracles of API Response Bodies for RESTful API Testing

@@ Abstract.
In API testing, deriving logical constraints on API response bodies to be used as oracles is crucial in generating test cases and performing automated testing of RESTful APIs. However, existing approaches are restricted to dynamic analysis in which oracles are extracted via the execution of APIs as part of the system under test. In this paper, we propose a complementary LLM-based, static approach in which the constraints for API response bodies are mined from API specifications. We leverage large language models (LLMs) to comprehend the API specifications, mine constraints for response bodies, and generate test cases. To reduce LLMs’ hallucination, we apply an Observation-Confirmation (OC) scheme which uses initial prompts to contextualize constraints, allowing subsequent prompts to more accurately confirm their presence. Our empirical results show that RBCTest with OC prompting achieves high precision in constraint mining with the average from 85.1%–93.6%. It also performs well in generating test cases from mined constraints, with a precision from 86.4%–91.7%. We also use the test cases generated by RBCTest to detect 46 mismatches between the API specification and actual response data for 19 real-world APIs. Four of the mismatches were, in fact, reported in developers’ forums.
AI4SE, Large Language Models, Automated Web Test Generation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; July 18, 2025; XXX, XXXjournalyear: 2026copyright: rightsretainedconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 1111111.1111111isbn: 1111111.111111ccs: Software and its engineering
@@ 1. Introduction
By adhering to the principles of Representational State Transfer (REST), the RESTful APIs provide a standardized way for interoperability among components and software systems. RESTful API testing helps identify and resolve several issues, ensuring that APIs perform as expected (viglianisi2020resttestgen; MartinLopez2021Restest; arcuri2019restful; liu2022morest; alonso2022arte; atlidakis2019restler). Among techniques for API testing, black-box testing uses the OpenAPI Specification (OAS) as a basis to generate test cases and data (kat; karlsson2020quickrest; MartinLopez2021Restest). The state-of-the-art API testing approaches are focused on status code (viglianisi2020resttestgen; kat; katalon; postman) and schema validation (viglianisi2020resttestgen; katalon; postman), even with rule extraction using human-readable descriptions in the OAS (kim2023enhancing). Existing static code validation approaches define an oracle for a test case by validating whether the response status code matches the expected value. In contrast, schema validation ensures the correctness of the response data by checking it against a specified schema.
@@
While status code and schema validation effectively cover aspects of data representation and status checking, they may overlook the logical correctness and validity of the response data from the APIs, which is essential. For example, an API request for a customer older than 18 receiving a response for one younger than 18 would not be detected by just validating the status code or schema. We also use the term oracles to denote such constraints on API response bodies.
@@
The state-of-the-art approaches for mining constraints on API response bodies focus only on dynamic analysis in which the constraints are extracted from the execution data of the system under test (SUT). AGORA/AGORA+ (alonso2023agora; Alonso2025AGORA_plus) automatically detects invariants—properties that should hold true. To identify invariants, it extends Daikon (ernst2007daikon), a dynamic instrumenter to detect invariants during the execution of the APIs within the systems.
@@
To complement such a dynamic approach as in AGORA/AGORA+, we present RBCTest, a novel LLM-based approach to mine the constraints of API response bodies from the API specification. To mine constraints, we leverage the ability of large language models (LLMs) to comprehend natural language descriptions in API specifications. Constraints on API response bodies are inferred from different sources, including response properties, response schema, operations, and request parameters. We also harness LLMs’ proficiency to create test cases from the mined constraints to verify if the SUT correctly returns the content satisfying the mined constraints. Moreover, we apply an Observation-Confirmation scheme by dividing the task of mining constraints into two phases: observation and confirmation. The initial prompt contextualizes the description of constraints, enabling the next prompt to more accurately decide their presence. As another issue in LLMs’ exploration capability, they could produce resulting constraints that are not true. Thus, we enhance RBCTest with two extra mechanisms. First, before requesting the LLM to make observations concerning constraints on parameters, we perform a filtering process to keep only the valid ones. Second, after generating test cases for mined constraints, we add a semantic verifier to verify those test cases against the examples specified in the OAS file. The idea is that such examples tend to be correct because they illustrate the descriptions on the data. For example, the OAS could give ”March” as a valid month. If such a correct example does not pass a test case generated by RBCTest based on the mined constraint(s), the test case must be incorrect, which is caused by incorrect constraint(s). Thus, our verifier will discard them, leading to improved precision.
@@
We evaluate RBCTest on two datasets, one from AGORA/AGORA+ (alonso2023agora) and our own dataset collected from 8 real-world API services consisting of 65 endpoints and 90 operations (website). Our dataset includes a benchmark containing manually verified constraints and test cases for response bodies of 8 SUTs, which serves as the ground-truth. Our empirical results show that RBCTest achieves high precision in constraint mining with the average of 93.6% for the RBCTest dataset and 85.1% for the AGORA/AGORA+ dataset. The precision values in test generation for constraint validation are from 86.4%–91.7%. We also leverage its generated tests to automatically verify mined constraints and detect mismatches between the API specifications and actual execution of the SUTs. A detected mismatch indicates a fault in the SUT or that the SUT does not match its specification. We report 46 actual faults found in 19 real-world applications, including 4 issues reported by users on GitLab Forum (gitlab; gitlab_issue_time1; gitlab_issue_time2; gitlab_issue_time3). This paper makes the following contributions:
@@
1. RBCTest: An LLM-based approach for constraint mining from API specifications and test generation for validating API response bodies.
@@
2. A manually-verified benchmark with ground-truth constraints for API response bodies of 8 real-world API services is available (website) for future research on API testing approaches.
3. An extensive evaluation showing RBCTest’s capabilities in constraint or oracle mining and test generation for API testing.
2. Motivation
Let us consider an example with Stripe, an online payment service. Figure 1(b) partially shows a description from the API specification (OAS), detailing the GET operation for retrieving charge records within a time interval (lines 9-10), which is defined by the gt and lt parameters (i.e., the lower and upper bounds). Users can also specify the customer for whom they wish to retrieve charging history (lines 15-17). A successful request returns a response with a status code of 200 and a response body with data. The structure of the response data (in Figure 1(a)) consists of a list of charge objects, each associated with various properties, e.g., amount, created timestamp, currency, and others. For instance, a GET request to the /v1/charges endpoint with parameters like ‘created[gt]=1679090500&customer=cus_idA’ returns a list of charges that satisfy the given conditions. An example response object is in Figure 2. The content of the response is constrained by the actual input parameters in the request. Thus, a complete testing process needs to verify the response content and the returned status code. For example, a test case can check if the created field of a returned charge object falls within the interval and ensuring that the customer field matches the requested ID.
@@ 2. Motivation
Observation 1 (Constraints from input parameters).
The response data is constrained by the input parameters from the request. When testing, in addition to verifying the status code, testers need to verify the response bodies returned from the APIs.
In Figure 1(a), the descriptions on the properties of the returned charge objects define certain constraints on the attributes. For example, the amount property must be positive, with a max value of eight digits. The currency attribute has a three-letter lowercase code.
Observation 2 (Constraints within response body).
Natural language descriptions on operations express logical constraints on operations, their responses, formatting requirements, or value range limitations that must be validated during API testing.
@@
3. Related Work
Recent surveys on API testing (kim2022automated; golmohammadi2022testing; ehsan2022restful; martin2022online; sharma2018automated; marculescu2022faults; martin2021black) reveal a trend towards automation adoption. AI/ML are used to enhance various aspects of API testing including of generating test cases (viglianisi2020resttestgen; MartinLopez2021Restest; atlidakis2019restler), realistic test inputs (alonso2022arte), and identify defects early in the development (liu2022morest; sahin2021discrete; arcuri2020automated; arcuri2021enhancing; arcuri2020handling; zhang2021adaptive; zhang2021enhancing; zhang2019resource; zhang2021resource). In API test case generation, validation typically falls into two main categories.
@@ 3. Related Work
Status Code Validation: Each HTTP request is returned with a response containing a status code and data. The status code, a 3-digit integer, indicates the outcome of the HTTP request. 2xx codes signify a successful request. Conversely, 4xx codes indicate errors, e.g., a bad syntax request or invalid input values. For instance, in testing the API ’GET/user_information’ with various valid and invalid user IDs, testers would expect the API to return status codes 200 or 404. This validation method is used in automated tools, e.g., Postman (postman), Katalon (katalon), RestTestGen (viglianisi2020resttestgen), and KAT (kat).
@@
Schema Validation: This verifies the presence of all required properties and the consistency of property data types with specifications. Tools, e.g., RestTestGen (viglianisi2020resttestgen), leverage external libraries, such as ‘swagger-schema-validator’, to facilitate schema validation.
@@
While status code and schema validation cover data representation and status checking, they overlook the logical correctness and validity in the API responses. For instance, if an API request for a charge in 2025 returns one from 2024, or if the amount is negative, these would not be detected by validating the status code or schema.
AGORA/AGORA+ (alonso2023agora; Alonso2025AGORA_plus) extends Daikon (ernst2007daikon) to infer the invariants from the execution data of the SUT using APIs. SATORI (alonso2025satori) uses LLMs to infer API’s expected behavior by analyzing the properties of the response fields of its operations. It does not have Observation-Confirmation (OC) scheme as in RBCTest.
KAT (kat) fully automates API testing using GPT and an input OpenAPI specification, building dependency graphs and generating test scripts. (kim2024leveraging) further applied GPT to enrich specifications with rule explanations and example inputs. Before LLMs, ARTE (alonso2022arte) used NLP to generate API test inputs. Morest (liu2022morest) proposed model-based RESTful API testing with a dynamic Property Graph, improving code coverage and bug detection. RESTler (atlidakis2019restler) introduced stateful fuzzing of REST APIs, inferring dependencies and generating tests guided by service responses. LlamaRestTest (kim2025LlamaRestTest) employs fine-tuned and quantized Llama3-8B model using mined datasets of REST API example values and inter-parameter dependencies, to generate realistic test inputs and uncover inter-parameter dependencies.
4. Key Ideas
@@
We design RBCTest, an LLM-based, static approach for mining the constraints of API response bodies and then generating test cases.
@@
Figure 3. Mining Constraints in APIs’ Response Bodies with LLMs
@@
Key Idea 1 [Leveraging LLMs to Mine Constraints in APIs’ response bodies]. RBCTest is a novel LLM-based static approach to mine the constraints for API’s response bodies from the OAS. The constraints on the response data can be found in the specification in either the descriptions of the operations (e.g., line 5 of Figure 1(a)) or the descriptions of the schema for the response data (lines 2–24 in Figure 1(b)). For example, the description in Figure 1 states that ‘the charges are returned in a sorted order with the most recent ones appearing first’. The schema for the returned values in Figure 1 also provides us several constraints on the parameters (lines 5–6) as well as the description of the returned object (line 2). For example, the amount of a charge is a positive number with up to 8 digits and such value must be of integer.
RBCTest is designed to complement dynamic approaches in scenarios where APIs and their execution is unavailable or unreliable. For example, this occurs when an API specification is available, but the APIs and their underlying systems are still under development.
@@
Key Idea 4 [Filter and Semantic Verifier]. Before asking the LLM to analyze constraints on parameters, we perform filtering to remove invalid constraints. After generating test cases based on the mined constraints, we introduce a semantic verifier to check these test cases against the examples in the OAS. The rationale is that these examples are typically accurate in the specification. For instance, an example of \$999,999.99 is used to illustrate a positive number for amount. If a test case fails to validate a given correct example, it suggests that the mined constraint may be incorrect.
5. Mining Constraints for API Response Bodies
The specification for an API endpoint includes two main parts: the request specification (the input of the API) and response schema specification (the output of the API). (1) A request specification determines how to call an API endpoint, detailing the required input parameters, their roles, and the parameters within the request body along with their respective roles. (2) The response schema specification provides instructions on the response body, including each property in the response data, its description, datatype, nullability, and other details. These parts are the target for constraint mining.